A projectile is launched with an initial velocity of 1.5km/s^-1 at an angle theta the target is located at a distance of 3000m and on the same level. The projectile strikes the target after 4.0s
a.determine the horizontal velocity of the projectile
b.calculate the value of theta

Grade:12th pass

1 Answers

Ravi
3 years ago
range=u2sin2\theta/2g
range=3000
u=1.5km/s is equal to 1500m/s
by above formula you will get \theta
then ucos\theta is the horizontal velocity